OpenStudy (anonymous):

Extracting the Roots m^2-75=0

OpenStudy (anonymous):

add 75 to both sides to get \[m^2=75\] then take the square roots. don't forget to write plus or minus, there are two answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

oh an one more thing \[\sqrt{75}=\sqrt{25\times 3}=\sqrt{25}\times \sqrt{3}=5\sqrt{3}\]